The length of a rectangle is 3cm longer than twice the width. the area of the rectangle is 90sqcm. Find the length and width of the rectangle
Answer by oberobic(2304)

L = 2W +3
.
L*W = 90
W = 90/L
.
substitute
.
L = 2(90/L) + 3
L = 180/L + 3
.
L^2 = 180 + 3L
.
L^2 -3L -180 = 0
.
(L-15)(L+12) = 0
.
L = 15 or L = -12
.
Negative length is not possible, so
L =15.
.
Substitute
.
W = 90/L
W = 90/15
W = 6
.
Check the area to be sure of the answer.
6*15 = 90
Correct.
.
Answer: Length = 15; width = 6.
.
Done.